# EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation

### v1

@@ Abstract
Summarization is a fundamental task in natural language processing (NLP) and since large language models (LLMs), such as GPT-4 and Claude, come out, increasing attention has been paid to long-form summarization whose input sequences are much longer, indicating more information contained. The current evaluation metrics either use similarity-based metrics like ROUGE and BERTScore which rely on similarity and fail to consider informativeness or LLM-based metrics, lacking quantitative analysis of information richness and are rather subjective. In this paper, we propose a new evaluation metric called EVA-Score using Atomic Fact Chain Generation and Document-level Relation Extraction together to automatically calculate the informativeness and give a definite number as an information score. Experiment results show that our metric shows a state-of-the-art correlation with humans. We also re-evaluate the performance of LLMs on long-form summarization comprehensively from the information aspect, forecasting future ways to use LLMs for long-form summarization.
EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation
Yuchen Fan1thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Chengsi Wang1, Gaochen Wu2, Bowen Zhou2thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Figure 1: Examples show the dilemma of ROUGE. The blue annotation is the atomic information contained in the text. Even though Candidate Summarization 1 is rewarded with a higher ROUGE Score, it contains less information than Candidate Summarization 2. It inspires us that ROUGE and BERTScore are not reasonable for evaluating summarization.
Summarization has long been a fundamental task in NLP and since LLMs came out (OpenAI et al., 2024; Touvron et al., 2023), many research leverages LLMs for summarization (Zhang et al., 2023b; Wang et al., 2023b). More recently, works on summarization moved to Long-form scenarios where the input sequence length is extremely larger, resulting in more information contained in it. Traditional evaluation metrics may be not suitable for evaluation of long-form summarization due to the great informativeness, thus calling for new evaluation metrics.
@@
Past evaluation metrics for summarization include some static evaluation methods, like ROUGE-k and ROUGE-l (Lin, 2004) which calculate the k-gram overlap or the longest common substring between the reference and the candidate. With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as a new evaluation metric based on pre-trained word embeddings. More recently, some works have begun to pay attention to evaluating long-form summarization (Kim et al., 2024; Chang et al., 2024) which uses prompts to define criteria to assign different scores for different aspects, like fluency, consistency, and coherence. However, traditional and LLM-based evaluation metrics face their dilemmas respectively.
@@
Figure 2: Pipeline of our evaluation metric. We begin with atomic fact generation. For generated facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement. Finally, when analyzing the truthfulness of a given fact, we use retrieval to find several reference facts and provide previous states.
For Similarity-based Metrics Since they are based on similarity, they cannot represent informativeness and are unsuitable in such an LLM era, where responses from LLMs are flexible and contain many synonyms and different syntax. For LLM-based Metrics They are insensitive to nuanced differences existing in summaries and their quantitative results are not reliable even if given detailed criteria for evaluation due to the black-box property.
We propose a new evaluation metric whose result is an objective and quantitative analysis but can fully catch the information contained in the summary. We call this metric Extraction-and-Validation Score. We begin with extracting atomic facts from reference and candidate summaries. Different from FActScore (Min et al., 2023), we claim that some information cannot live without a context, so we treat them as part of logic chains. Afterward, we introduce Document-level Relation Extraction to remedy the loss of hierarchical relations. Then for each atomic fact to be verified, we use information retrieval to find the most relevant atomic facts as references and prompt LLMs to focus only on newly added information and decide whether it is entailed or not. Intuitively, LLMs for evaluation can be more accurate in single verification than in complex situations. We ensure that only one piece of information is evaluated at a time, simplifying the validation process. And by introducing Document-level Relations, we form a more comprehensive information set, maintaining the completeness of the evaluation.
We conduct comprehensive experiments by comparing EVA-Score with current evaluation metrics, including traditional and LLM-based. The results show that EVA-Score achieves the best human alignment compared to SOTA past evaluation metric, BERTScore by 0.166, 0.120, and 0.09 in Pearson, Spearman, and Kendall correlation respectively at the text level. At the system level, it shows human-comparable evaluation, with 0.975 in Pearson correlation and 1.0 in both Spearman and Kendall correlation. We also re-evaluate long-form summarization based on EVA-Score and the experiment results show the informativeness of different LLMs on long-form summarization. The results show the great performance of GPT-4 when the context becomes longer, ranking first in arXiv and Gov Report and second in Pubmed and BookSum. It also shows the promising performance of Mixtral-7B-Instruct for long-form summarization.
Our main contributions are as follows:
  • We propose a new evaluation metric that can generate high-quality evaluation results of a given summary, based on atomic fact chain generation and document-level relation extraction, by providing a numerical score and steps to obtain this score.
  • We conduct comprehensive experiments on several long-form summarization datasets, with the number of tokens of each record ranging from 1k to 14k. Experiment results show that our evaluation metric shows superiority in every correlation with humans.
  • We re-evaluate long-form summarization concerning informativeness which is the first to evaluate from the information aspect as far as we know. In our experiments, the state-of-the-art model GPT-4 still performs better than any other models in long context settings but lags behind some assistant models in short context settings.
@@ 2 Related Work
2.1 Long-form Summarization
Long-form summarization requires language models to understand the purpose of the context and capture core information from a long context which is a struggle for language models. During the last few years, much work from different fields regarding long-form summarization has been studied such as news reports (See et al., 2017), government reports (Cohan et al., 2018), academic papers and books (Kryściński et al., 2022), posing great challenges for language models (Lewis et al., 2019; Raffel et al., 2023). Since ChatGPT and GPT-4 (OpenAI et al., 2024) come out, LLMs show great potential to deal with long-form summarization (Lu et al., 2023; Zhang et al., 2023a; Wang et al., 2023a) because of their huge amount of pre-training corpus. However, when the context length becomes much larger, LLMs still have trouble because of limitations for input sequences and knowledge forgetting when reading a long context. More recently, some research aims at exploring LLMs which can handle longer context paving the way for long-form summarization. (Su et al., 2023; Li* et al., 2023) Yet, summarizing report-level or book-level texts is still a difficult task. (Chang et al., 2024)
@@ 2.2 Evaluation of Summarization
Tradition evaluation metrics for summarization include ROUGE variants, like ROUGE-k and ROUGE-L and BERTScore (Zhang et al., 2020). However, they only reflect the similarity between candidates and references without the information contained in the prediction which is not enough for us to choose a better summary or quantize the informativeness. Currently, several evaluation metrics (Zheng et al., 2024; Ke et al., 2023) based on LLMs have been proposed devoted to aligning model evaluation with human evaluation. Some works apply LLMs to the evaluation of long-form summarization, for example, Luo et al. (2023) uses ChatGPT to evaluate factual inconsistency for text summarization. More Recently, (Jung et al., 2024) calculates the saliency of a summary using MSM(Masked Language Model) to find how a candidate summary can be recovered given a reference summary. However, we claim that making use of the inherent ability of LLMs only is not enough to measure the ability of models to summarize for the lack of evaluation of the degree of information contained. Moreover, leveraging pre-trained probability distribution is not good enough because of its unexplainability. We hope to find a reasonable and explainable method for evaluation.
@@ 3 Preliminaries
In this section, we provide definitions for elements used in our evaluation pipeline.
Our work begins with extracting atomic facts from documents. Specifically, we denote reference summary as $s$ and candidate summary as $\hat{s}$. The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being single atomic facts in the list. After atomic fact chain generation, we obtain lists containing logic chains $C$ and $\hat{C}$ for reference and candidate, while the elements of them are defined as $C_{i}$ and $\hat{C_{i}}$ and atomic fact in logic chains are defined as $C_{i_{j}}$ and ${\hat{C_{i_{j}}}}$. Finally, we conduct Document-level Relation Extraction(DocRE) to complement the original atomic fact list with hierarchical information. The final lists are $DocS$ and $Doc\hat{S}$ separately.
@@ 4 Methods
In this section, we introduce our proposed pipeline to evaluate informativeness in long-form summarization by extraction and validation. Given $s$ and $\hat{s}$, our automatic evaluation metric operates by 1) obtaining all atomic facts from $s$ and $\hat{s}$; 2) reforming atomic facts to form logic chains; 3) extracting all document-level relations and filter the duplicated one; 4) using LLM to validate whether it is true or false.
Figure 3: Examples show the generation of atomic facts. It uses demonstration to help LLMs understand the task.
@@ 4.1 Atomic Fact Generation
Atomic facts are defined as concise sentences that convey a single piece of information. In any given sentence within a model prediction, there may be multiple pieces of information that require individual verification. Consequently, evaluating the truthfulness of sentences on a one-by-one basis renders the true-or-false outcomes inadequate for assessing informativeness. Previous studies find that splitting candidate prediction into atomic facts (Min et al., 2023) improves the evaluation by a large margin and enables a more fine-grained evaluation. Following their work, we use ChatGPT with same prompt to generate atomic facts from both $S$ and $\hat{S}$ as $\left[s_{1},s_{2},s_{3}...\right]$ and $\left[\hat{s_{1}},\hat{s_{2}},\hat{s_{3}}...\right]$ where $s_{i}$ is atomic fact from reference summarization and $\hat{s_{i}}$ is atomic fact from candidate summarization.
4.2 Atomic Fact Chain Generation
Although claimed atomic, an examination of the “atomic facts” presented reveals that they often encompass more than one atomic fact. As illustrated in Figure 3, certain information cannot be independently extracted due to its reliance on contextual semantics. Consequently, in cases of cascaded atomic facts, where subsequent facts incorporate or entirely encompass information from preceding facts, if any of the preceding facts are incorrect, all subsequent facts are deemed incorrect, even if the newly added information is accurate.
To clarify, consider the examples: He played for Berekum Arsenal” and He played for Berekum Arsenal before moving to the United States.” These are two cascaded facts, with the latter not being atomic. However, the latter sentence cannot be decomposed into atomic facts because the adverbial clause depends on the entirety of the sentence for its meaning. Consequently, if the initial statement is incorrect, any subsequent statements, including those with additional correct information, are also considered incorrect.
We propose using LLMs, e.g. Mixtral-7B-instruct in our work to find logical relationships between atomic facts and treating atomic facts as a list of logic chains. The reformed lists are represented as $[C_{1},C_{2},C_{3}...]$ and $[\hat{C_{1}},\hat{C_{2}},\hat{C_{3}}...]$.
@@ 4.3 Document-level Relation Extraction
Extracting atomic facts from text primarily captures sentence-level relations. However, it is crucial to also consider document-level relations, which are challenging to discern solely through atomic fact generation. These relations are integral to understanding the contextual logic of a document. We complement Document-level relation extractions(DocRE) using the state-of-the-art model GPT-4 with in-context learning for both $s$ and $\hat{s}$ and the results are $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2})...]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{t_{2}},\hat{r_{2}})...]$ where $(h_{i},r_{i},t_{i})$ and $(\hat{h_{i}},\hat{r_{i}},\hat{t_{i}})$ are head entity, relation and tail entity correspondingly. We then paraphrase the triples to make them natural language for post-processing.
To ensure that no sentence-level relations are included and contaminate work later, we use BERTScore along with TF-IDF to filter out the similar ones existing both in the atomic fact list and document-level relation list.
@@ 4.4 LLM for Validation
For each atomic fact $s_{i}$ in the candidate atomic fact list, we retrieve $k$ most relevant atomic facts from the reference summarization atomic fact list where k is a hyper-parameter which is a trade-off between accuracy and efficiency. For each atomic fact $\hat{C_{i_{j}}}$ in the corresponding logic chain $\hat{C_{i}}$, we check out one by one according to $s_{i}$ from $k$ most relevant atomic facts in $DocS$ represented as $f(\hat{C_{i_{j}}},s_{i})$. Specifically, we use Mixtral-7B-Instruct as our validation model. By providing the validation model with previous states and prompting it not to pay attention to previous information, we ensure that one piece of information is treated at a time.
@@
$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{C_{i_{j}}},s_{i})\right)$ (1)
$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(s_{i},\hat{C_{i_{j}}})\right)$ (2)

### v3

# EVA-Score: Evaluating Abstractive Long-form Summarization on Informativeness through Extraction and Validation

@@ Abstract
Since LLMs emerged, more attention has been paid to abstractive long-form summarization, where longer input sequences indicate more information contained. Nevertheless, the automatic evaluation of such summaries remains underexplored. Current evaluation metrics either use traditional metrics like ROUGE and BERTScore, which rely on surface-level similarity and fail to consider informativeness, or simple LLM-based metrics, which are not robust and easily overwhelmed by the long contexts. In this paper, we propose a new evaluation metric called EVA-Score to extract all information from the given summaries, identify overlapped information based on reference, and calculate the information score. We test EVA-Score on several datasets and the experimental results reveal that EVA-Score shows the highest correlation with humans. We also re-evaluate the performance of LLMs on long-form summarization from the information perspective. The results indicate that responses of LLMs still have a gap with the human-written answers. Moreover, we provide a detailed analysis of the effectiveness of EVA-Score, forecasting future ways to automatically evaluate abstractive long-form summarization.
EVA-Score: Evaluating Abstractive Long-form Summarization on Informativeness through Extraction and Validation
Yuchen Fan1,thanks:  The work was done when he worked as an intern at Tsinghua University, China., Xin Zhong1, Yazhe Wan1, Chengsi Wang1, Haonan Cheng1, Gaochen Wu2, Ning Ding2, Bowen Zhou2,thanks:  Corresponding authors 1 Beijing University of Posts and Telecommunications 2 Department of Electronic Engineering, Tsinghua University yuchenfan48@gmail.com zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Figure 1: The example shows the dilemma of traditional evaluation metrics, including ROUGE and BERTScore. Candidate Summary 1 and Candidate Summary 2 are summaries to be evaluated. The information within the summary is denoted using blue. Even though Candidate Summary 1 is rewarded with a higher ROUGE and BERTScore, it contains less information and is given a lower EVA-Score which meets human preference. It inspires us that traditional metrics are not reasonable for evaluating long-form and knowledge-intensive summaries due to their insensitivity to information.
Long-form summarization is challenging due to its extensive knowledge and lengthy context. It can be categorized into extractive (Zhang et al., 2023) and abstractive (Wang et al., 2023b), where the former retains exact content from the document, while the latter produces a paraphrased and coherent output. With advancements in LLMs (OpenAI et al., 2024; Touvron et al., 2023), significant progress has been made (Chang et al., 2024; Chhikara et al., 2024; Kim et al., 2024). However, the automatic evaluation of abstractive long-form summarization remains underexplored and is more challenging than evaluating extractive ones due to the implicit equivalence of information involved. Therefore, an information-sensitive metric is needed to effectively capture the depth and breadth of summaries.
@@
Past evaluation metrics for abstractive summarization include some static evaluation methods, like ROUGE-k and ROUGE-L (Lin, 2004). With the development of pre-trained language models, BERTScore (Zhang et al., 2020) is widely used as an evaluation metric based on word embeddings to measure the similarity between the reference and the candidate. Also, some works use LLMs to evaluate long-form summarization, defining criteria, and prompting LLMs to assign scores for different aspects like fluency and consistency. For clarity, we categorize metrics that utilize prompts and obtain responses from LLMs as LLM-based evaluation metrics (Kim et al., 2024; Chang et al., 2024), while other metrics like ROUGE and BERTScore are referred to as traditional metrics. Each category, however, has its limitations and challenges, especially when applied to complex, nuanced content in abstractive long-form summarization.
@@
Figure 2: Pipeline of EVA-Score. We begin with Atomic Fact Generation. For extracted facts, we use LLMs to check whether they can be treated as a logic chain and use Document-level Relation Extraction as a complement. “Doc Re” in the figure is actually “Document-level Relation” for abstraction. When analyzing the truthfulness of a given fact, we retrieve and find several reference facts and provide the evaluation results of previous facts within a logic chain for help. Thus, we ensure that one piece of information is analyzed once.
Traditional metrics, such as ROUGE and BERTScore, are primarily based on surface-level similarity—either word-level or embedding-level similarity to the reference (Krishna et al., 2021). Therefore, they often fail to capture information-level similarity, which involves understanding the actual meaning or key ideas conveyed, making them less suitable in the era of LLMs, where responses are highly flexible and often use synonyms or different syntactic structures. We show an example in Figure 1, indicating the inherent drawback of ROUGE and BERTScore. On the other hand, LLM-based Metrics struggle with capturing nuanced information differences between candidates and references. Also, their evaluations can be unreliable, even with detailed evaluation criteria in such a complex scenario, where LLMs must first identify key information from the candidate and then assess it against the reference (Shen et al., 2023).
To fill the gap, we propose a new evaluation metric called Extraction-and-Validation Score (EVA-Score), designed to provide an objective and quantitative analysis while capturing the full range of information in a summary. Our approach begins by extracting atomic facts from reference and candidate. Unlike FActScore (Min et al., 2023), we recognize that certain pieces of information are context-dependent, resulting in some extracted facts not being atomic but cascaded between them instead. To address this, we reformat these atomic facts into logic chains that preserve contextual integrity. Additionally, since atomic fact extraction operates on a sentence-level granularity, it often misses document-level relationships. For complement, we incorporate Document-Level Relation Extraction (Xue et al., 2024; Xu et al., 2022) to capture the hierarchical relations and provide a more complete understanding of the content. To verify the accuracy of each atomic fact, we identify the most relevant context from the reference and prompt LLMs to determine whether the candidate information is entailed by the reference. In our implementation, we ensure that only one piece of information is evaluated at a time, which simplifies the validation and enhances reliability. The overall evaluation pipeline is illustrated in Figure 2.
To the best of our knowledge, we are the first to evaluate long-form summarization from the informativeness perspective. We conduct comprehensive experiments to compare EVA-Score with existing evaluation metrics, including both traditional and LLM-based metrics. The results reveal that EVA-Score achieves the highest alignment with human evaluations, significantly outperforming past metrics. At the system level, EVA-Score demonstrates a strong correlation with human judgments, achieving a Pearson correlation of 0.975 and perfect scores of 1.0 in both Spearman and Kendall correlations. Additionally, we re-evaluate long-form summarization performance based on EVA-Score. The results highlight the impressive capability of GPT-4, especially as the context length increases, where it ranks first on two test datasets and second on the remaining two. We also analyze why certain LLMs fail on specific datasets and provide insights into scenarios where EVA-Score aligns with human annotations versus cases where discrepancies occur. These analyses shed light on the future development of automatic evaluation metrics for abstractive long-form summarization.
@@ 2 Related Work
2.1 Abstractive Long-form Summarization
With the advent of models like ChatGPT and GPT-4 (OpenAI et al., 2024), LLMs have demonstrated considerable potential in traditional summarization tasks (Lu et al., 2023; Zhang et al., 2023; Wang et al., 2023a). However, as context length grows, LLMs still face significant issues (Bai et al., 2023). Recent research has explored methods to extend the capability of LLMs to handle longer contexts, which meanwhile facilitates long-form summarization (Su et al., 2023; Li* et al., 2023). Long-form summarization requires language models to understand the purpose of the context and paraphrases core information from extensive text which can be categorized into extractive and abstractive summarization, with the latter providing a more fluent and flexible summary that better aligns with human preferences compared to extractive methods. Therefore, many studies focus on abstractive long-form summarization in different domains, including news reports (See et al., 2017), government reports (Cohan et al., 2018), and books (Kryściński et al., 2022). Despite these efforts, abstractive summarization for book-level and report-level texts still remains a challenging task (Chang et al., 2024).
@@ 2.2 Evaluation of Summarization
Tradition evaluation metrics for summarization include ROUGE and its variants, and BERTScore (Zhang et al., 2020). However, (Krishna et al., 2021) highlights that ROUGE is not such an informative metric, implying its drawback for abstractive long-form summarization. Similar challenges apply to BERTScore. Currently, several LLM-based evaluation metrics (Zheng et al., 2024; Ke et al., 2023) have been proposed to aligning model evaluation with human evaluation. Some works apply LLMs to the evaluation of long-form summarization, for example, Luo et al. (2023) uses ChatGPT to evaluate factual inconsistency for text summarization and Song et al. (2024) leverages LLMs to evaluate summaries based on faithfulness, completeness, and conciseness. However, we argue that only using the inherent ability of LLMs is not enough to measure the quality of a summary due to the inferior ability of LLMs to differentiate between summaries with nuanced information. We are seeking to find a reasonable yet explainable method for the evaluation of long-form summarization.
@@ 3 Preliminaries
In this section, we provide definitions for terms used in our evaluation pipeline.
Our work begins with extracting atomic facts from the reference and the candidate which we denote as $Sum$ and as $\hat{Sum}$ respectively. The extracted atomic fact lists are $S$ and $\hat{S}$ with $s_{i}$ and $\hat{s_{i}}$ being each atomic fact in the list. After Atomic Fact Chain Generation, we obtain fact chain lists $\hat{C}$ from $\hat{S}$, while each element in it is defined $\hat{C_{i}}$ and each atomic fact in logic chains is defined as ${\hat{C_{i_{j}}}}$. Finally, we conduct Document-level Relation Extraction to complement the original atomic fact list with hierarchical information. The final lists are $DocS$ and $Doc\hat{S}$ separately and the elements in them are denoted as $D_{i}$ and ${\hat{D_{i_{j}}}}$.
@@ 4 Methods
In this section, we introduce the detailed implementation of EVA-Score to evaluate informativeness in long-form summarization.
@@ 4.1 Atomic Fact Generation
In any given sentence within a model prediction, multiple pieces of information may coexist. Consequently, evaluating sentences as a single unit renders the true-or-false outcomes, which are inadequate for assessing the quality. Min et al. (2023) finds that splitting the prediction into atomic facts, defined as concise sentences consisting of a single piece of information, enables a more fine-grained evaluation and improves the evaluation by a large margin. Following their work, we use ChatGPT with the same prompt to generate atomic facts from both $Sum$ and $\hat{Sum}$ and obtain $S$ and $\hat{S}$. We share an example in Appendix A.1. The prompt used for Atomic Fact Generation is displayed in Appendix C.1. The temperature is set to $0.0$ to ensure the effectiveness of demonstrations.
4.2 Atomic Fact Chain Generation (AFCG)
Although claimed atomic, certain information, like time and location, cannot be independently extracted due to its reliance on contextual semantics. Consequently, parts of $\hat{S}$ may appear as cascaded atomic fact lists, where subsequent facts incorporate or entirely encompass information from preceding facts. In this case, if any of the preceding facts are incorrect, all subsequent facts are deemed wrong, even if the newly added information is accurate. Since we try to measure the quality of a summary from the information perspective, the former incorrect information should not affect the latter. To mitigate the impact, we reform the atomic facts into several logic chains so that in one logic chain, when analyzing a certain atomic fact, we can explicitly focus on newly added information. We show an example of AFCG in Appendix A.2. In particular, AFCG applies only to $\hat{S}$ since all the information in $Sum$ is considered ground truth. Therefore, the cascade effect does not exist.
We use Mistral-7B-instruct (Jiang et al., 2023a) in our work to find logical relationships between atomic facts, treating it as an NLI task between one fact and the exact next fact, due to the special structure of $\hat{S}$ (Min et al., 2023), and reforming atomic facts as a list of logic chains. The prompt used for AFCG is listed in Appendix C.2. The temperature is set at $1.0$ to encourage more divergent responses.
@@ 4.3 Document-level Relation Extraction
Extracting atomic facts from text mainly captures sentence-level relations—those between entities within the same sentence. However, document-level relations, involving entities across an entire document or beyond sentence boundaries, are crucial for understanding contextual logic. To address this, we implement Document-level Relation Extraction (DocRE) using GPT-4, starting with Named Entity Recognition (NER) followed by in-context learning to extract document-level relations from $Sum$ and $\hat{Sum}$. The outputs are tuples $[(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots]$ and $[(\hat{h_{1}},\hat{r_{1}},\hat{t_{1}}),(\hat{h_{2}},\hat{r_{2}},\hat{t_{2}}),\dots]$, representing head entities, relations, and tail entities respectively. The temperature for NER and DocRE is set at $1.0$ to ensure diverse outputs. Appendix A.3 provides an example illustrating differences between sentence-level and document-level relations, and Appendices 14 and C.4 show the prompts for NER and DocRE, respectively. Finally, the triples are paraphrased into natural language for post-processing.
To prevent the inclusion of sentence-level relations, we use cosine similarity based on BERT (Devlin et al., 2019) to filter out overlapping information between the atomic fact list and the document-level relation list. To determine the threshold, we randomly sample a subset from the datasets listed in Section 5.1, test various thresholds, and select the optimal one, as detailed in Appendix B.2. We ultimately choose a threshold of $0.65$. The filtered relations are then incorporated into $C$ and $\hat{C}$ to form $DocS$ and $\hat{DocS}$ where $DocS$ includes atomic facts from Section 4.1 and document-level relations, whereas $\hat{DocS}$ comprises atomic fact chains from Section 4.2 and document-level relations with each relation as a single-element list.
@@ 4.4 LLM for Validation
For each atomic fact $\hat{D_{i_{j}}}$ in $\hat{DocS}$, we retrieve $k$ most relevant atomic facts according to BERTScore (Zhang et al., 2020) from $DocS$, using $\hat{D_{i_{j}}}$ as hypothesis and each $D_{i}$ as the reference. $k$ is a hyperparameter which is a trade-off between effectiveness and efficiency, and we use $k=3$ in our experiment settings. We use Mistral-7B-instruct as our validation model with a temperature equal to $1.0$. The matching process between two statements is represented using $f(\cdot)$. By providing the correctness of previous statements and prompting it not to pay attention to them but to focus on newly added information, we ensure that one piece of information is checked each time. We provide an example in Appendix A.4 and the prompt used for LLMs is listed in Appendix C.5.
@@
$\text{Precision}=\frac{1}{|Doc\hat{S}|}\sum_{i=0}^{n}\mathds{1}\left(f(\hat{D_{i_{j}}},D_{i})\right)$ (1)
$\text{Recall}=\frac{1}{|DocS|}\sum_{i=0}^{n}\mathds{1}\left(f^{-1}(D_{i},\hat{D_{i_{j}}})\right)$ (2)